# Neuronal correlations shape the scaling behavior of memory capacity and nonlinear computational capability of recurrent neural networks

### v2

@@ A.2 Statistics for neuronal correlations
for $i\neq j$. From the ansatz, $[\langle\phi_{i}\phi_{j}\rangle^{2}]_{(i\neq j)}\sim O(1/N)$, we see that $[\langle x_{i}x_{j}\rangle^{2}]_{(i\neq j)}\sim O(1/N)$.
Consequently, similar to single-neuron statistics, the non-trivial perturbation induced by auxiliary neurons, i.e., the second and third terms in Eq.(A.2), does not contribute to the leading order of neuronal correlations. In contrast, in the case of a continuous-time RNN, the perturbation induced by auxiliary neurons contributes to the leading order of neuronal correlations [20]. This discrepancy stems from the difference in the strength of the autocorrelation, $\langle x_{i}(t)x_{i}(t-\tau)\rangle_{t}$, which scales as $O(1)$ in the continuous-time model, but as $O(1/\sqrt{N})$ in our discrete-time model. This property of negligible perturbation effects by auxiliary neurons in a discrete-time model greatly simplifies the analytical derivation of memory capacity (see Appendix.B.1). Note that even in a discrete-time model, perturbation-induced effects contribute to the leading-order term of neuronal correlations if the RNN dynamics include a leak term, i.e. $x_{i}(t+1)=\gamma x_{i}(t)+\sum_{j}^{N}J_{ij}\phi_{j}(t)$ with $\gamma>0$, which yields an $O(1)$ autocorrelation.
To calculate $[\langle\phi_{i}\phi_{j}\rangle^{2}]$, we introduce Price’s theorem [19]. For two Gaussian variables, $\begin{pmatrix}z_{1}\\ z_{2}\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}c_{1}&\tau\\ \tau&c_{2}\\ \end{pmatrix}\right)$, and any smooth activation function $\phi$, we define $f_{\phi}(c_{1},c_{2},\tau)\equiv\langle\phi(z_{1})\phi(z_{2})\rangle$. Price’s theorem then states:
@@
which is self-consistent with the ansatz, $[\langle\phi_{i}\phi_{j}\rangle^{2}]\sim O(1/N)$.
A.3 Proof of $10<g\langle\phi^{\prime}(x)\rangle_{*}<$ 1
We prove an inequality, $10<g\langle\phi^{\prime}(x)\rangle_{*}<$ 1 for any $g>0$, $\sigma_{s}^{2}\geq 0$, $\sigma_{n}^{2}\geq 0$, and an odd saturated sigmoid function satisfying $\phi^{\prime}(0)=1$, $\phi^{\prime}(x)>0$, $\phi^{\prime\prime}(x)\leq 0\ (x\geq 0)$ and $\phi(\pm\infty)=\pm 1$. We use the abbreviation, $K\equiv[\langle x_{i}^{2}\rangle]$. It is evident that $g\langle\phi^{\prime}(x)\rangle_{*}>0$ because $g>0$ and $\phi^{\prime}(x)\geq 0$. We prove that $g\langle\phi^{\prime}(x)\rangle_{*}<1$ below.
@@ B.1 Analytical derivation of memory capacity
$C^{-1}=\sum_{n=0}^{\infty}D^{-1}\left(-\tilde{C}D^{-1}\right)^{n},$ (48)
where $D_{ij}\equiv\delta_{ij}\langle x_{i}(t)^{2}\rangle$ and $\tilde{C}_{ij}\equiv(1-\delta_{ij})\langle x_{i}(t)x_{j}(t)\rangle$.
This series expansion is ensured to converge when the spectrum norm is less than one, $\|\tilde{C}D^{-1}\|<1$. However, calculating the exact value of the spectrum norm is a challenging task. Employing the inequality $\|\tilde{C}D^{-1}\|\leq\|\tilde{C}D^{-1}\|_{F}$, where $\|\cdot\|_{F}$ denotes a Frobenius norm, we can obtain the sufficient condition for the convergence, $\|\tilde{C}D^{-1}\|_{F}<1$. The value of $\|\tilde{C}D^{-1}\|_{F}$ is calculated as
$\|\tilde{C}D^{-1}\|_{F}$ $=\sum_{i,j(i\neq j)}^{\alpha\sqrt{N}}\frac{\langle x_{i}x_{j}\rangle^{2}}{\langle x_{i}^{2}\rangle^{2}}=\alpha^{2}N\left[\frac{\langle x_{i}x_{j}\rangle^{2}_{(i\neq j)}}{\langle x_{i}^{2}\rangle^{2}}\right]=\alpha^{2}N\frac{[\langle x_{i}x_{j}\rangle^{2}]_{(i\neq j)}}{[\langle x_{i}^{2}\rangle]^{2}}=\frac{\alpha^{2}}{[\langle x_{i}^{2}\rangle]^{2}}\frac{g^{4}\langle\phi(x)^{2}\rangle_{*}^{2}+\tilde{\sigma}_{s}^{4}}{1-g^{4}\langle\phi^{\prime}(x)\rangle_{*}^{4}},$ (49)
@@
For the solution of memory capacity, it is enough to calculate each quenched averaging term in Eq.(B.1). We perform this calculation by the dynamical cavity method, similar to the approach described in Appendix.A. As noted there, for a reservoir RNN with discrete time dynamics, perturbations induced by the auxiliary neurons do not contribute to the leading order term, which simplifies the calculation.
To begin with, we calculate the first term. Employing Eq.(28) and introducing an auxiliary neuron indexed by 00, it is enough to calculate $[\langle s_{t-d}x_{0,t}\rangle^{2}]$ as:
@@
where we used the independence between $J_{0,i},J_{0,j}$ and $\phi_{i,t}$ in the third equation. Since $(s_{t},x_{i,t})$ are Gaussian variables, we apply Price’s theorem, resulting in
@@
where we restored $[\langle s_{t-(d-1)}x_{i,t}\rangle^{2}]$ to $[\langle s_{t-(d-1)}x_{0,t}\rangle^{2}]$. This equation is a recurrence formula for $[\langle s_{t-d}x_{0,t}\rangle^{2}]$, whose solution is
@@
Next, we move on to the calculation of the second term. We introduce the two auxiliary neurons indexed by 00 and $0^{\prime}$, obtaining for $d\geq 1$
@@
Subsequently, we proceed to calculate the third term in Eq.(B.1). It is enough to consider the case for $i\neq k$ because the terms for $i=k$ have the same order as those for $i\neq k$, but the number of $i=k$ terms, $O(N^{2})$, is much smaller than those of $i\neq k$ terms, $O(N^{3})$, making their overall contribution subleading. We introduce three auxiliary neurons, labeled 00, $0^{\prime}$, and $0^{\prime\prime}$, obtaining for $d\geq 1$
@@
B.2 Analytical solution of the decay rate, $\lim_{N\to\infty}r(L)$
@@
Here, in the first inequality, we apply the result $10<g\langle\phi^{\prime}(x)\rangle_{*}<$ 1 proven in Appendix.A.3, and $\frac{\alpha\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}<1$ as shown below:
@@
B.3 Decay rate for a linear RNN
@@
[…]
@@ C.1 Overview of IPC theory
$C_{T}[f]\equiv 1-\frac{{\rm{min}}_{\bm{w}}\langle(z(t)-f(t))^{2}\rangle_{T}}{\langle f(t)^{2}\rangle_{T}},$ (73)
where $\langle\cdot\rangle_{T}$ denotes the time average over simulation time $T$. It is proven that $0\leq C_{T}[f]\leq 1$ holds true for any function $f$. The total IPC is subsequently defined as the summation of capacities across all basis polynomials:
$IPC_{\rm{total}}\equiv\sum_{\bm{d}}C_{T}[y_{\bm{d}}].$ (74)
It has been established that $\lim_{T\to\infty}IPC_{\rm{total}}$ is equivalent to the number of readout units, $L$, provided that the reservoir satisfies echo state property [2], whereby the reservoir’s state is uniquely determined solely by input signals, regardless of its initial state.
@@
$IPC_{D}\equiv\sum_{\begin{subarray}{c}\bm{d}\\ {\rm{s.t.}}\ deg(y_{\bm{d}})=D\end{subarray}}C_{T}[y_{\bm{d}}].$ (75)
Crucially, $IPC_{1}$ is identical to memory capacity, as $C_{T}[y_{\bm{d}}]$ for $deg(y_{\bm{d}})=1$ corresponds to Eq.(2). In contrast, $IPC_{D}$ for $D\geq 2$ represents a non-linear computational ability of the reservoir.
When calculating the capacities by numerical simulation, we must take care not to overestimate them, because for finite simulation time $T$, they are subject to a systematic positive error. Following the approach of Dambre et al. [7], we determine the threshold of the capacities, $\epsilon$, as
@@
where $\chi^{2}(L)$ denotes a random variable that follows a Chi-squared distribution with $L$ degrees of freedom. The value of $p$ represents the probability that a truly zero capacity is incorrectly assessed as non-zero. We assume $C_{T}=0$ if $C_{T}$ is lower than a threshold value $\epsilon$. Following Dambre et al., we set $p=10^{-4}$ for all our numerical simulations of memory capacity ($=IPC_{1}$) and IPC, although the choice of $p$ has a negligible impact on our results.
@@
C.2 Simulation results for IPC for an RNN with the ReLU activation function
Our theoretical framework for the analytical derivation of memory capacity cannot be applied to the RNNs with the ReLU activation function. This limitation arises because Price’s theorem (Eq.(37)) cannot be used for the ReLU function, as it is non-smooth at the origin. Therefore, we numerically obtain memory capacity and IPC of a ReLU RNN, as shown in Fig.8. Similarly to the RNNs employing the sigmoid function, the memory capacity ($IPC_{1}$) increases sublinearly with $L$, while non-linear computational capacities ($IPC_{D}$ for $D\geq 2$) emerge supralinearly and sequentially.
Note that unlike the sigmoid RNNs, the ReLU RNNs do not exhibit chaotic behavior. Instead, they diverge when the parameter $g$ exceeds a critical value, $g_{c}$. To demonstrate this, substituting $\phi(x)={\rm max}(0,x)$ into the dynamical mean-field equation (Appendix.A.1), we obtain
$[x_{i}(t+1)^{2}]=\frac{g^{2}}{2}[x_{i}(t)^{2}]+s(t+1)^{2}+\sigma_{n}^{2}+O(1/\sqrt{N}).$ (78)
This self-consistent equation admits a stable solution only when $g<g_{c}=\sqrt{2}$. For $g>g_{c}$, the system diverges regardless of the parameters $\sigma_{s}$ and $\sigma_{n}$. Consequently, the numerical simulations presented in Fig.8 are conducted with $g$ values strictly less than $\sqrt{2}$.
Figure 8: IPC values for varying the number of readout neurons $L$ of a reservoir RNN with the ReLU activation function $\phi(x)={\rm max}(0,x)$ and network size $N=1000$. The figure’s structure and content mirror Fig.5. The values of IPC for up to degree 5 are calculated. Simulation time length is $T=10^{4}$. The model parameters $(g,\sigma_{s},\sigma_{n})$ are $(1.0,1.0,0.0)$ for (a) and $(1.0,1.0,0.3)$ for (b). The RNN for (a) satisfies the echo state property since its positive maximum conditional Lyapunov exponent is negtive ($\lambda_{\rm max}=-0.35\pm 0.007$), while the property breaks down for (b) as the RNN receives neuronal noise.
C.3 Linear scaling of $IPC_{D}$ with $L$ for vanishing neuronal correlations
In the main text, we claim that each $IPC_{D}$ for $D\geq 1$ exhibits linear scaling with L if we ignore neuronal correlations. The following provides a concise proof of this claim. First, similarly to Eq.(4), the capacity for a basis function $y_{\bm{d}}$ can be expressed as

### v3

@@ A.2 Statistics for neuronal correlations
for $i\neq j$. From the ansatz, $\langle\phi_{i}\phi_{j}\rangle^{2}]_{(i\neq j)}\sim O(1/N)$, we see that $\langle x_{i}x_{j}\rangle^{2}]_{(i\neq j)}\sim O(1/N)$.
Consequently, similar to single-neuron statistics, the non-trivial perturbation induced by auxiliary neurons, i.e., the second and third terms in Eq.(A.2), does not contribute to the leading order of neuronal correlations. In contrast, in the case of a continuous-time RNN, the perturbation induced by auxiliary neurons contributes to the leading order of neuronal correlations [26]. This discrepancy stems from the difference in the strength of the autocorrelation, $\langle x_{i}(t)x_{i}(t-\tau)\rangle_{t}$, which scales as $O(1)$ in the continuous-time model, but as $O(1/N^{\tau/2})$ in our discrete-time model. This property of negligible perturbation effects by auxiliary neurons in a discrete-time model greatly simplifies the analytical derivation of memory capacity (see Appendix B.1). Note that even in a discrete-time model, perturbation-induced effects contribute to the leading-order term of neuronal correlations if the RNN dynamics include a leak term, i.e. $x_{i}(t+1)=\gamma x_{i}(t)+\sum_{j}^{N}J_{ij}\phi_{j}(t)$ with $\gamma>0$, which yields an $O(1)$ autocorrelation.
To calculate $\langle\phi_{i}\phi_{j}\rangle^{2}]$, we introduce Price’s theorem [25]. For two Gaussian variables, $\begin{pmatrix}z_{1}\\ z_{2}\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}c_{1}&\tau\\ \tau&c_{2}\\ \end{pmatrix}\right)$, and any smooth activation function $\phi$, we define $f_{\phi}(c_{1},c_{2},\tau)\equiv\langle\phi(z_{1})\phi(z_{2})\rangle$. Price’s theorem then states:
@@
which is self-consistent with the ansatz, $\langle\phi_{i}\phi_{j}\rangle^{2}]\sim O(1/N)$.
A.3 Proof of $0<g\langle\phi^{\prime}(x)\rangle_{*}<1$
We prove an inequality, $0<g\langle\phi^{\prime}(x)\rangle_{*}<1$ for any $g>0$, $\sigma_{s}^{2}\geq 0$, $\sigma_{n}^{2}\geq 0$, and an odd saturated sigmoid function satisfying $\phi^{\prime}(0)=1$, $\phi^{\prime}(x)>0$, $\phi^{\prime\prime}(x)\leq 0\ (x\geq 0)$ and $\phi(\pm\infty)=\pm 1$. We use the abbreviation, $K\equiv[\langle x_{i}^{2}\rangle]$. It is evident that $g\langle\phi^{\prime}(x)\rangle_{*}>0$ because $g>0$ and $\phi^{\prime}(x)\geq 0$. We prove that $g\langle\phi^{\prime}(x)\rangle_{*}<1$ below.
@@ B.1 Analytical derivation of memory capacity
$C^{-1}=\sum_{n=0}^{\infty}C_{\rm diag}^{-1}\left(-C_{\rm nondiag}C_{\rm diag}^{-1}\right)^{n},$ (53)
where $(C_{\rm diag})_{ij}\equiv\delta_{ij}\langle x_{i}(t)^{2}\rangle$ and $(C_{\rm nondiag})_{ij}\equiv(1-\delta_{ij})\langle x_{i}(t)x_{j}(t)\rangle$.
This series expansion is ensured to converge when the spectral norm is less than one, $\|C_{\rm nondiag}C_{\rm diag}^{-1}\|<1$. However, calculating the exact value of the spectral norm is a challenging task. Employing the inequality $\|C_{\rm nondiag}C_{\rm diag}^{-1}\|\leq\|C_{\rm nondiag}C_{\rm diag}^{-1}\|_{F}$, where $\|\cdot\|_{F}$ denotes a Frobenius norm, we can obtain the sufficient condition for the convergence, $\|C_{\rm nondiag}C_{\rm diag}^{-1}\|_{F}<1$. The value of $\|C_{\rm nondiag}C_{\rm diag}^{-1}\|_{F}$ is calculated as
$\|C_{\rm nondiag}C_{\rm diag}^{-1}\|_{F}$ $=\sum_{i,j(i\neq j)}^{\alpha\sqrt{N}}\frac{\langle x_{i}x_{j}\rangle^{2}}{\langle x_{i}^{2}\rangle^{2}}=\alpha^{2}N\left[\frac{\langle x_{i}x_{j}\rangle^{2}_{(i\neq j)}}{\langle x_{i}^{2}\rangle^{2}}\right]=\alpha^{2}N\frac{[\langle x_{i}x_{j}\rangle^{2}]_{(i\neq j)}}{[\langle x_{i}^{2}\rangle]^{2}}=\frac{\alpha^{2}}{[\langle x_{i}^{2}\rangle]^{2}}\frac{g^{4}\langle\phi(x)^{2}\rangle_{*}^{2}+\tilde{\sigma}_{s}^{4}}{1-g^{4}\langle\phi^{\prime}(x)\rangle_{*}^{4}},$ (54)
@@
For the solution of memory capacity, it is enough to calculate each quenched averaging term in Eq.(B.1). We perform this calculation by the dynamical cavity method, similar to the approach described in appendix A. As noted there, for a reservoir RNN with discrete time dynamics, perturbations induced by the auxiliary neurons do not contribute to the leading order term, which simplifies the calculation.
To begin with, we calculate the first term. Employing Eq.(31) and introducing an auxiliary neuron indexed by 0, it is enough to calculate $\langle s_{t-d}x_{0,t}\rangle^{2}]$ as:
@@
where we used the independence between $J_{0,i},J_{0,j}$ and $\phi_{i,t}$ in the third equation. To proceed, we need to evaluate $\langle s_{t-d}\phi(x_{i,t})\rangle^{2}]$. Since $(s_{t-d},x_{i,t})$ are correlated Gaussian variables, we can express them using auxiliary variables $y,z$:
$s_{t-d}$ $=\sqrt{\sigma_{s}^{2}-\frac{\langle s_{t-d}x_{0,t}\rangle^{2}}{\langle x_{0}^{2}\rangle}}\ y+\frac{\langle s_{t-d}x_{0,t}\rangle}{\sqrt{\langle x_{0}^{2}\rangle}}z$
$x_{0,t}$ $=\sqrt{\langle x_{0}^{2}\rangle}\ z.$ (59)
Using these expressions, we can calculate $\langle s_{t-d}\phi(x_{i,t})\rangle$ as
$\langle s_{t-d}\phi(x_{i,t})\rangle$ $=\int Dy\ Dz\ \left(\sqrt{\sigma_{s}^{2}-\frac{\langle s_{t-d}x_{0,t}\rangle^{2}}{\langle x_{0}^{2}\rangle}}\ y+\frac{\langle s_{t-d}x_{0,t}\rangle}{\sqrt{\langle x_{0}^{2}\rangle}}z\right)\phi\left(\sqrt{\langle x_{0}^{2}\rangle}\ z\right)$ (60)
$=\frac{\langle s_{t-d}x_{0,t}\rangle}{\sqrt{\langle x_{0}^{2}\rangle}}\int Dz\ z\ \phi\left(\sqrt{\langle x_{0}^{2}\rangle}\ z\right)$
$=\langle s_{t-d}x_{0,t}\rangle\int Dz\ \phi^{\prime}\left(\sqrt{\langle x_{0}^{2}\rangle}\ z\right)$
$=\langle s_{t-d}x_{0,t}\rangle\langle\phi^{\prime}(x_{0})\rangle,$
where $Dy$ and $Dz$ denotes standard Gaussian measures, and integration by parts is employed to derive the third line. Therefore, we obtain
@@
where we restored $\langle s_{t-(d-1)}x_{i,t}\rangle^{2}]$ to $\langle s_{t-(d-1)}x_{0,t}\rangle^{2}]$. This equation is a recurrence formula for $\langle s_{t-d}x_{0,t}\rangle^{2}]$, whose solution is
@@
Next, we move on to the calculation of the second term. We introduce the two auxiliary neurons indexed by 0 and $0^{\prime}$, obtaining for $d\geq 1$
@@
Subsequently, we proceed to calculate the third term in Eq.(B.1). It is enough to consider the case for $i\neq k$ because the terms for $i=k$ have the same order as those for $i\neq k$, but the number of $i=k$ terms, $O(N^{2})$, is much smaller than those of $i\neq k$ terms, $O(N^{3})$, making their overall contribution subleading. We introduce three auxiliary neurons, labeled 0, $0^{\prime}$, and $0^{\prime\prime}$, obtaining for $d\geq 1$
@@
B.2 Generalization of scaling assumptions
In this subsection, our aim is to identify broader scaling assumptions – beyond those specified in Eqs.(4) – under which our theoretical framework for analytically deriving memory capacity remains applicable. To this end, we consider the following generalized scaling relations:
$L=\alpha N^{c_{1}},\ \ \sigma_{s}^{2}=\frac{\tilde{\sigma}_{s}^{2}}{N^{c_{2}}},\ \ \sigma_{n}^{2}\sim O(1),$ (73)
where $c_{1},c_{2}\geq 0$.
First, we require that the memory capacity remains non-vanishing in the limit $N\to\infty$. The first-order term in the Neumann series expansion of $M_{d}]$ is given by
$\frac{1}{\sigma_{s}^{2}}\sum_{i}^{\alpha N^{c_{1}}}\left[\frac{\langle s_{t-d}x_{i,t}\rangle^{2}}{\langle x_{i}^{2}\rangle}\right]\sim O(N^{c_{1}-c_{2}}).$ (74)
To ensure that the memory capacity does not vanish as $N\to\infty$, it is necessary that $c_{1}\geq c_{2}$.
Second, from the sufficient condition for the convergence of the Neumann series expansion, we obtain an additional constraint on $c_{1}$ and $c_{2}$. Specifically, the Frobenius norm of the matrix $C_{\rm nondiag}C_{\rm diag}^{-1}$ must satisfy
$\|C_{\rm nondiag}C_{\rm diag}^{-1}\|_{F}=\alpha^{2}N^{2c_{1}}\frac{[\langle x_{i}x_{j}\rangle^{2}]_{(i\neq j)}}{[\langle x_{i}^{2}\rangle]^{2}}\sim O(1).$ (75)
The order of the neuronal correlations $\langle x_{i}x_{j}\rangle^{2}]_{(i\neq j)}$ can be obtained using the same calculation as described in appendix A.2, yielding $\langle x_{i}x_{j}\rangle^{2}]_{(i\neq j)}\sim O(1/N^{\min(1,2c_{2})})$. Substituting this result into Eq.(75), we obtain
$\|C_{\rm nondiag}C_{\rm diag}^{-1}\|_{F}\sim O(N^{2c_{1}-\min({1,2c_{2}})}),$ (76)
from which it follows that $2c_{1}\leq\min(1,2c_{2})$ must be satisfied.
To summarize, the scaling exponents, $c_{1}$ and $d_{2}$, must satisfy the following two conditions:
$c_{1}$ $\geq c_{2}$ (77)
$2c_{1}$ $\leq\min(1,2c_{2}),$
Combining these inequalities, we conclude that the allowed values are $0\leq c_{1}=c_{2}\leq\frac{1}{2}$.
B.3 Analytical solution of the decay rate, $\lim_{N\to\infty}r(L)$
@@
Here, in the first inequality, we apply the result $0<g\langle\phi^{\prime}(x)\rangle_{*}<1$ proven in appendix A.3, and $\frac{\alpha\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}<1$ as shown below:
@@
B.4 Decay rate for a linear RNN
@@
Figure 9: (a) Memory capacity and (b) growth rate of memory capacity for linear RNNs. The figure’s structure and content are the same as Fig.2, with analytical and numerical results presented in the same format. The input intensity is $\sigma_{s}^{2}=1.0^{2}/\sqrt{N}$ ($\tilde{\sigma}_{s}^{2}=1.0^{2}$). In simulations, the network size is $N=10000$, and simulation time length is $T=10^{4}$.
@@ C.1 Overview of IPC theory
$C[f]\equiv 1-\frac{{\rm{min}}_{\bm{w}}\langle(z(t)-f(t))^{2}\rangle_{t}}{\langle f(t)^{2}\rangle_{t}},$ (85)
where $\langle\cdot\rangle_{t}$ denotes the time average. It is proven that $0\leq C[f]\leq 1$ holds true for any function $f$. The total IPC is subsequently defined as the summation of capacities across all basis polynomials:
$IPC_{\rm{total}}\equiv\sum_{\bm{d}}C[y_{\bm{d}}].$ (86)
Notably, it has been established that $IPC_{\rm total}=L$ holds, provided that the reservoir’s activity is full-rank and the reservoir satisfies the fading memory property [13], whereby the reservoir’s state is uniquely determined solely by input signals, regardless of its initial state.
@@
$IPC_{D}\equiv\sum_{\begin{subarray}{c}\bm{d}\\ {\rm{s.t.}}\ deg(y_{\bm{d}})=D\end{subarray}}C[y_{\bm{d}}].$ (87)
Crucially, $IPC_{1}$ is identical to memory capacity, as $C[y_{\bm{d}}]$ for $deg(y_{\bm{d}})=1$ corresponds to Eq.(3). In contrast, $IPC_{D}$ for $D\geq 2$ represents a non-linear computational ability of the reservoir.
When calculating the capacities by numerical simulation, we must take care not to overestimate them, because for a finite simulation time $T$, they are subject to a systematic positive error. Following Dambre et al. [13], we determine the threshold of the capacities, $\epsilon$, as
@@
where $\chi^{2}(L)$ denotes a random variable that follows a Chi-squared distribution with $L$ degrees of freedom. The value of $p$ represents the probability that a truly zero capacity is incorrectly assessed as non-zero. We consider a capacity to be zero if it falls below this threshold $\epsilon$. Consistent with Dambre et al., we set $p=10^{-4}$ for all our numerical simulations of memory capacity ($IPC_{1}$) and IPC, although the choice of $p$ has a negligible impact on our results.
@@
C.2 Linear scaling of $IPC_{D}$ with $L$ for vanishing neuronal correlations
In the main text, we claim that each $IPC_{D}$ for $D\geq 1$ exhibits linear scaling with regard to $L$ if we ignore neuronal correlations. The following provides a concise proof of this claim. First, similarly to Eq.(5), the capacity for a basis function $y_{\bm{d}}$ can be expressed as